OpenStudy (anonymous):

write it in the form of a+ib (1+i)2(1-i)3 1+i square and 1-i cube

OpenStudy (watchmath):

\((1+i)^2(1-i)^2(1-i)=2^2(1-i)=4-4i\)

OpenStudy (anonymous):

1+i is square and 1-i is cube

OpenStudy (watchmath):

Yes, but we can write \((1+i)^3\) as \((1+i)^2(1+i)\) right ?

OpenStudy (watchmath):

I mean \((1-i)^3=(1-i)^2(1-i)\)

OpenStudy (anonymous):

my answer comes-4+4i

OpenStudy (watchmath):

How did you get that?

OpenStudy (anonymous):

thanx a lot for the help:)